\begin{thebibliography}{36}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Aharoni et~al.(2019)Aharoni, Johnson, and
  Firat}]{aharoni-etal-2019-massively}
Roee Aharoni, Melvin Johnson, and Orhan Firat. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1388} {Massively multilingual
  neural machine translation}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 3874--3884,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Anastasopoulos et~al.(2020)Anastasopoulos, Cattelan, Dou, Federico,
  Federmann, Genzel, Guzm{\'a}n, Hu, Hughes, Koehn, Lazar, Lewis, Neubig, Niu,
  {\"O}ktem, Paquin, Tang, and Tur}]{anastasopoulos-etal-2020-tico}
Antonios Anastasopoulos, Alessandro Cattelan, Zi-Yi Dou, Marcello Federico,
  Christian Federmann, Dmitriy Genzel, Franscisco Guzm{\'a}n, Junjie Hu,
  Macduff Hughes, Philipp Koehn, Rosie Lazar, Will Lewis, Graham Neubig,
  Mengmeng Niu, Alp {\"O}ktem, Eric Paquin, Grace Tang, and Sylwia Tur. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.nlpcovid19-2.5} {{TICO}-19:
  the translation initiative for {CO}vid-19}.
\newblock In \emph{Proceedings of the 1st Workshop on {NLP} for {COVID}-19
  (Part 2) at {EMNLP} 2020}, Online. Association for Computational Linguistics.

\bibitem[{Arivazhagan et~al.(2019)Arivazhagan, Bapna, Firat, Lepikhin, Johnson,
  Krikun, Chen, Cao, Foster, Cherry, Macherey, Chen, and
  Wu}]{https://doi.org/10.48550/arxiv.1907.05019}
Naveen Arivazhagan, Ankur Bapna, Orhan Firat, Dmitry Lepikhin, Melvin Johnson,
  Maxim Krikun, Mia~Xu Chen, Yuan Cao, George Foster, Colin Cherry, Wolfgang
  Macherey, Zhifeng Chen, and Yonghui Wu. 2019.
\newblock \href {https://doi.org/10.48550/ARXIV.1907.05019} {Massively
  multilingual neural machine translation in the wild: Findings and
  challenges}.

\bibitem[{Behnke and Heafield(2020)}]{behnke-heafield-2020-losing}
Maximiliana Behnke and Kenneth Heafield. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.211} {Losing heads
  in the lottery: Pruning transformer attention in neural machine translation}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 2664--2674, Online. Association
  for Computational Linguistics.

\bibitem[{Berard et~al.(2021)Berard, Lee, Clinchant, Jung, and
  Nikoulina}]{berard-etal-2021-efficient}
Alexandre Berard, Dain Lee, Stephane Clinchant, Kweonwoo Jung, and Vassilina
  Nikoulina. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.674} {Efficient
  inference for multilingual neural machine translation}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 8563--8583, Online and Punta Cana,
  Dominican Republic. Association for Computational Linguistics.

\bibitem[{Dong et~al.(2015)Dong, Wu, He, Yu, and Wang}]{dong-etal-2015-multi}
Daxiang Dong, Hua Wu, Wei He, Dianhai Yu, and Haifeng Wang. 2015.
\newblock \href {https://doi.org/10.3115/v1/P15-1166} {Multi-task learning for
  multiple language translation}.
\newblock In \emph{Proceedings of the 53rd Annual Meeting of the Association
  for Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 1723--1732,
  Beijing, China. Association for Computational Linguistics.

\bibitem[{Edunov et~al.(2018)Edunov, Ott, Auli, and
  Grangier}]{edunov-etal-2018-understanding}
Sergey Edunov, Myle Ott, Michael Auli, and David Grangier. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1045} {Understanding
  back-translation at scale}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 489--500, Brussels, Belgium. Association
  for Computational Linguistics.

\bibitem[{El-Kishky et~al.(2020)El-Kishky, Chaudhary, Guzm{\'a}n, and
  Koehn}]{el-kishky-etal-2020-ccaligned}
Ahmed El-Kishky, Vishrav Chaudhary, Francisco Guzm{\'a}n, and Philipp Koehn.
  2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.480} {{CCA}ligned:
  A massive collection of cross-lingual web-document pairs}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 5960--5969, Online. Association
  for Computational Linguistics.

\bibitem[{Fan et~al.(2020)Fan, Bhosale, Schwenk, Ma, El-Kishky, Goyal, Baines,
  Celebi, Wenzek, Chaudhary, Goyal, Birch, Liptchinsky, Edunov, Grave, Auli,
  and Joulin}]{m2m-100}
Angela Fan, Shruti Bhosale, Holger Schwenk, Zhiyi Ma, Ahmed El-Kishky,
  Siddharth Goyal, Mandeep Baines, Onur Celebi, Guillaume Wenzek, Vishrav
  Chaudhary, Naman Goyal, Tom Birch, Vitaliy Liptchinsky, Sergey Edunov,
  Edouard Grave, Michael Auli, and Armand Joulin. 2020.
\newblock \href {https://doi.org/10.48550/ARXIV.2010.11125} {Beyond
  english-centric multilingual machine translation}.

\bibitem[{Firat et~al.(2016)Firat, Cho, and Bengio}]{firat-etal-2016-multi}
Orhan Firat, Kyunghyun Cho, and Yoshua Bengio. 2016.
\newblock \href {https://doi.org/10.18653/v1/N16-1101} {Multi-way, multilingual
  neural machine translation with a shared attention mechanism}.
\newblock In \emph{Proceedings of the 2016 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 866--875, San Diego, California. Association for
  Computational Linguistics.

\bibitem[{Garcia et~al.(2021)Garcia, Siddhant, Firat, and
  Parikh}]{garcia-etal-2021-harnessing}
Xavier Garcia, Aditya Siddhant, Orhan Firat, and Ankur Parikh. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.89} {Harnessing
  multilinguality in unsupervised machine translation for rare languages}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 1126--1137, Online. Association for Computational
  Linguistics.

\bibitem[{Goyal et~al.(2021)Goyal, Gao, Chaudhary, Chen, Wenzek, Ju, Krishnan,
  Ranzato, Guzman, and Fan}]{flores101}
Naman Goyal, Cynthia Gao, Vishrav Chaudhary, Peng-Jen Chen, Guillaume Wenzek,
  Da~Ju, Sanjana Krishnan, Marc'Aurelio Ranzato, Francisco Guzman, and Angela
  Fan. 2021.
\newblock \href {https://doi.org/10.48550/ARXIV.2106.03193} {The flores-101
  evaluation benchmark for low-resource and multilingual machine translation}.

\bibitem[{Haddow et~al.(2021)Haddow, Bawden, Barone, Helcl, and
  Birch}]{haddow2021survey}
Barry Haddow, Rachel Bawden, Antonio Valerio~Miceli Barone, Jindřich Helcl,
  and Alexandra Birch. 2021.
\newblock \href {http://arxiv.org/abs/2109.00486} {Survey of low-resource
  machine translation}.

\bibitem[{Hu et~al.(2018)Hu, Peng, Wei, Huang, Li, Yang, and
  Zhou}]{hu-etal-2018-attention}
Minghao Hu, Yuxing Peng, Furu Wei, Zhen Huang, Dongsheng Li, Nan Yang, and Ming
  Zhou. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1232} {Attention-guided answer
  distillation for machine reading comprehension}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 2077--2086, Brussels, Belgium.
  Association for Computational Linguistics.

\bibitem[{Johnson et~al.(2017)Johnson, Schuster, Le, Krikun, Wu, Chen, Thorat,
  Vi{\'e}gas, Wattenberg, Corrado, Hughes, and
  Dean}]{johnson-etal-2017-googles}
Melvin Johnson, Mike Schuster, Quoc~V. Le, Maxim Krikun, Yonghui Wu, Zhifeng
  Chen, Nikhil Thorat, Fernanda Vi{\'e}gas, Martin Wattenberg, Greg Corrado,
  Macduff Hughes, and Jeffrey Dean. 2017.
\newblock \href {https://doi.org/10.1162/tacl_a_00065} {{G}oogle{'}s
  multilingual neural machine translation system: Enabling zero-shot
  translation}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  5:339--351.

\bibitem[{Kasai et~al.(2021)Kasai, Pappas, Peng, Cross, and
  Smith}]{kasai2021_deepshallow}
Jungo Kasai, Nikolaos Pappas, Hao Peng, James Cross, and Noah Smith. 2021.
\newblock \href {https://openreview.net/forum?id=KpfasTaLUpq} {Deep encoder,
  shallow decoder: Reevaluating non-autoregressive machine translation}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Kim and Rush(2016)}]{kim-rush-2016-sequence}
Yoon Kim and Alexander~M. Rush. 2016.
\newblock \href {https://doi.org/10.18653/v1/D16-1139} {Sequence-level
  knowledge distillation}.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pages 1317--1327, Austin, Texas. Association
  for Computational Linguistics.

\bibitem[{Ko et~al.(2021)Ko, El-Kishky, Renduchintala, Chaudhary, Goyal,
  Guzm{\'a}n, Fung, Koehn, and Diab}]{ko-etal-2021-adapting}
Wei-Jen Ko, Ahmed El-Kishky, Adithya Renduchintala, Vishrav Chaudhary, Naman
  Goyal, Francisco Guzm{\'a}n, Pascale Fung, Philipp Koehn, and Mona Diab.
  2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.66} {Adapting
  high-resource {NMT} models to translate low-resource related languages
  without parallel data}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 802--812, Online.
  Association for Computational Linguistics.

\bibitem[{Li et~al.(2021)Li, Wang, Liu, Du, Xiao, Zhang, and Zhu}]{light2021}
Bei Li, Ziyang Wang, Hui Liu, Quan Du, Tong Xiao, Chunliang Zhang, and Jingbo
  Zhu. 2021.
\newblock \href {https://doi.org/10.1609/aaai.v35i15.17561} {Learning
  light-weight translation models from deep transformer}.
\newblock \emph{Proceedings of the AAAI Conference on Artificial Intelligence},
  35(15):13217–13225.

\bibitem[{Mohammadshahi et~al.(2022)Mohammadshahi, Nikoulina, Berard, Brun,
  Henderson, and Besacier}]{mohammadshahi2022compressed}
Alireza Mohammadshahi, Vassilina Nikoulina, Alexandre Berard, Caroline Brun,
  James Henderson, and Laurent Besacier. 2022.
\newblock \href {http://arxiv.org/abs/2205.10828} {What do compressed
  multilingual machine translation models forget?}

\bibitem[{Neubig and Hu(2018)}]{neubig-hu-2018-rapid}
Graham Neubig and Junjie Hu. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1103} {Rapid adaptation of
  neural machine translation to new languages}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 875--880, Brussels, Belgium. Association
  for Computational Linguistics.

\bibitem[{Platanios et~al.(2018)Platanios, Sachan, Neubig, and
  Mitchell}]{platanios-etal-2018-contextual}
Emmanouil~Antonios Platanios, Mrinmaya Sachan, Graham Neubig, and Tom Mitchell.
  2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1039} {Contextual parameter
  generation for universal neural machine translation}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 425--435, Brussels, Belgium. Association
  for Computational Linguistics.

\bibitem[{Raffel et~al.(2019)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{raffel2019exploring}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael
  Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2019.
\newblock \href {http://arxiv.org/abs/1910.10683} {Exploring the limits of
  transfer learning with a unified text-to-text transformer}.

\bibitem[{Schwenk et~al.(2019)Schwenk, Wenzek, Edunov, Grave, and
  Joulin}]{https://doi.org/10.48550/arxiv.1911.04944}
Holger Schwenk, Guillaume Wenzek, Sergey Edunov, Edouard Grave, and Armand
  Joulin. 2019.
\newblock \href {https://doi.org/10.48550/ARXIV.1911.04944} {Ccmatrix: Mining
  billions of high-quality parallel sentences on the web}.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch}]{sennrich-etal-2016-improving}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock \href {https://doi.org/10.18653/v1/P16-1009} {Improving neural
  machine translation models with monolingual data}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 86--96, Berlin,
  Germany. Association for Computational Linguistics.

\bibitem[{Tang et~al.(2021)Tang, Tran, Li, Chen, Goyal, Chaudhary, Gu, and
  Fan}]{tang-etal-2021-multilingual}
Yuqing Tang, Chau Tran, Xian Li, Peng-Jen Chen, Naman Goyal, Vishrav Chaudhary,
  Jiatao Gu, and Angela Fan. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.findings-acl.304}
  {Multilingual translation from denoising pre-training}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL-IJCNLP 2021}, pages 3450--3466, Online. Association for Computational
  Linguistics.

\bibitem[{Tao et~al.(2022)Tao, Hou, Zhang, Shang, Jiang, Liu, Luo, and
  Wong}]{tao2022compression}
Chaofan Tao, Lu~Hou, Wei Zhang, Lifeng Shang, Xin Jiang, Qun Liu, Ping Luo, and
  Ngai Wong. 2022.
\newblock \href {http://arxiv.org/abs/2203.10705} {Compression of generative
  pre-trained language models via quantization}.

\bibitem[{Tiedemann(2012)}]{tiedemann-2012-parallel}
J{\"o}rg Tiedemann. 2012.
\newblock \href
  {http://www.lrec-conf.org/proceedings/lrec2012/pdf/463_Paper.pdf} {Parallel
  data, tools and interfaces in {OPUS}}.
\newblock In \emph{Proceedings of the Eighth International Conference on
  Language Resources and Evaluation ({LREC}-2012)}, pages 2214--2218, Istanbul,
  Turkey. European Languages Resources Association (ELRA).

\bibitem[{Tiedemann(2020)}]{tiedemann-2020-tatoeba}
J{\"o}rg Tiedemann. 2020.
\newblock \href {https://aclanthology.org/2020.wmt-1.139} {The tatoeba
  translation challenge {--} realistic data sets for low resource and
  multilingual {MT}}.
\newblock In \emph{Proceedings of the Fifth Conference on Machine Translation},
  pages 1174--1182, Online. Association for Computational Linguistics.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{transformer}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, \L~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href
  {https://proceedings.neurips.cc/paper/2017/file/3f5ee243547dee91fbd053c1c4a845aa-Paper.pdf}
  {Attention is all you need}.
\newblock In \emph{Advances in Neural Information Processing Systems},
  volume~30. Curran Associates, Inc.

\bibitem[{Wang et~al.(2021)Wang, Yan, Meng, and
  Zhou}]{wang-etal-2021-selective}
Fusheng Wang, Jianhao Yan, Fandong Meng, and Jie Zhou. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.504} {Selective
  knowledge distillation for neural machine translation}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 6456--6466,
  Online. Association for Computational Linguistics.

\bibitem[{Wu et~al.(2020)Wu, Passban, Rezagholizadeh, and
  Liu}]{wu-etal-2020-skip}
Yimeng Wu, Peyman Passban, Mehdi Rezagholizadeh, and Qun Liu. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.74} {Why skip if
  you can combine: A simple knowledge distillation technique for intermediate
  layers}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 1016--1021, Online. Association
  for Computational Linguistics.

\bibitem[{Yao et~al.(2022)Yao, Aminabadi, Zhang, Wu, Li, and
  He}]{yao2022zeroquant}
Zhewei Yao, Reza~Yazdani Aminabadi, Minjia Zhang, Xiaoxia Wu, Conglong Li, and
  Yuxiong He. 2022.
\newblock \href {http://arxiv.org/abs/2206.01861} {Zeroquant: Efficient and
  affordable post-training quantization for large-scale transformers}.

\bibitem[{Zhang et~al.(2020)Zhang, Williams, Titov, and
  Sennrich}]{zhang-etal-2020-improving}
Biao Zhang, Philip Williams, Ivan Titov, and Rico Sennrich. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.148} {Improving
  massively multilingual neural machine translation and zero-shot translation}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 1628--1639, Online. Association for
  Computational Linguistics.

\bibitem[{Zhang et~al.(2022)Zhang, Roller, Goyal, Artetxe, Chen, Chen, Dewan,
  Diab, Li, Lin, Mihaylov, Ott, Shleifer, Shuster, Simig, Koura, Sridhar, Wang,
  and Zettlemoyer}]{zhang2022opt}
Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui
  Chen, Christopher Dewan, Mona Diab, Xian Li, Xi~Victoria Lin, Todor Mihaylov,
  Myle Ott, Sam Shleifer, Kurt Shuster, Daniel Simig, Punit~Singh Koura, Anjali
  Sridhar, Tianlu Wang, and Luke Zettlemoyer. 2022.
\newblock \href {http://arxiv.org/abs/2205.01068} {Opt: Open pre-trained
  transformer language models}.

\bibitem[{Zhang et~al.(2021)Zhang, Huang, Feng, and
  Cao}]{zhang-etal-2021-enlivening}
Tianfu Zhang, Heyan Huang, Chong Feng, and Longbing Cao. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.260} {Enlivening
  redundant heads in multi-head self-attention for machine translation}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 3238--3248, Online and Punta Cana,
  Dominican Republic. Association for Computational Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{EMNLP2022}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage[utf8]{inputenc}

\usepackage{microtype}
\usepackage{rotating}
\usepackage{blindtext}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{mathtools}
\usepackage{booktabs}
\usepackage{relsize}
\usepackage{float}
\usepackage{graphics}
\usepackage{multirow}
\usepackage{algorithm}
\usepackage{algpseudocode}
\usepackage[utf8]{inputenc}

\usepackage{threeparttable}
\usepackage{booktabs, caption, makecell}

\usepackage[title]{appendix}
\usepackage[cal=cm]{mathalfa}
\usepackage[colorinlistoftodos]{todonotes}
\usepackage{multirow}
\usepackage[export]{adjustbox}
\usepackage{subcaption}
\usepackage{pifont}
\usepackage{tabularx}
\usepackage{longtable}
\usepackage{tikz}
\usepackage{enumitem}
\usepackage{diagbox}
\usepackage{bbm}

\usepackage{colortbl,dcolumn}
\usepackage{microtype}
\usepackage[subtle]{savetrees}
\title{SMaLL-100: Introducing \textbf{S}hallow Multilingual Machine Translation \textbf{M}odel for \textbf{L}ow-Resource \textbf{L}anguages}

\author{Alireza Mohammadshahi\thanks{~~~Work done during an internship at NAVER LABS Europe.}~~$^{1,2,3}$ ~~Vassilina Nikoulina$^1$ ~~Alexandre Berard$^1$ \\ \textbf{Caroline Brun}$^1$ ~~\textbf{James Henderson}$^2$ ~~\textbf{Laurent Besacier}$^1$ \vspace{0.1cm}\\ 
 $^1$ NAVER LABS Europe ~~~~~ $^2$ IDIAP Research Institute ~~~~~ $^3$ EPFL\\
 \texttt{\{first.last\}@naverlabs.com}\\\texttt{\{alireza.mohammadshahi,james.henderson\}@idiap.ch}
}

\begin{document}
\maketitle

\begin{abstract}

In recent years,  multilingual machine translation models have achieved promising performance on low-resource language pairs by sharing information between similar languages, thus enabling zero-shot translation.  To overcome the "curse of multilinguality", these models often opt for scaling up the number of parameters, which makes their use in resource-constrained environments challenging. We introduce \textit{SMaLL-100}, a distilled version of the M2M-100~(12B) model, a massively multilingual machine translation model covering 100 languages. We train SMaLL-100 with uniform sampling across all language pairs and therefore focus on preserving the performance of low-resource languages.  We evaluate SMaLL-100 on different low-resource benchmarks: FLORES-101, Tatoeba, and TICO-19 and demonstrate that it outperforms previous massively multilingual models of comparable sizes (200-600M) while improving inference latency and memory usage. Additionally, our model achieves comparable results to M2M-100~(1.2B), while being 3.6$\times$ smaller and 4.3$\times$ faster at inference.\footnote{The code and pre-trained SMaLL-100 model is available at \url{https://github.com/alirezamshi/small100}.}

\end{abstract}

\section{Introduction}

Neural Machine Translation~(NMT) systems are usually trained on datasets consisting of millions of parallel sentences, thus still performing poorly on low-resource languages, i.e., languages without a large amount of training data. 
Over the past few years, previous work has proposed several approaches to improve the quality of translations in low-resource languages, e.g., Multilingual Neural Machine Translation~(MNMT) models~\cite{johnson-etal-2017-googles,m2m-100,tang-etal-2021-multilingual,flores101}, back-translation~\cite{sennrich-etal-2016-improving,edunov-etal-2018-understanding} and unsupervised machine translation~\cite{garcia-etal-2021-harnessing,ko-etal-2021-adapting}. 
Massively MNMT models are particularly interesting for low-resource languages as they benefit the most from knowledge transfer from related languages~\cite{https://doi.org/10.48550/arxiv.1907.05019}. However, it is also seen that \textit{curse of multilinguality} hurts the performance of high-resource languages. So, previous work attempted to increase the model size to maintain the translation performance in both high and low-resource languages. This makes the use of these massively MNMT models challenging in real-world resource-constrained environments. To overcome this problem, we propose SMaLL-100, a \textbf{S}hallow \textbf{M}ultilingual M\textbf{a}chine Translation Model for \textbf{L}ow-Resource \textbf{L}anguages covering 100 languages, which is a distilled alternative of M2M-100~(12B)~\cite{m2m-100}, the most recent and biggest available multilingual NMT model. In this paper, we focus on very-low and low-resource language pairs as there is no reasonable-size universal model that achieves acceptable performance over a great number of low-resource languages. We do so by training SMaLL-100 on a perfectly balanced dataset.\footnote{All language pairs have the same sampling probability, regardless of their training data size.} While this leads to lower performance on the high-resource languages, we claim that this loss is easily recoverable through further fine-tuning.  We evaluate SMaLL-100 on different low-resource benchmarks, e.g., FLORES-101~\cite{flores101}, Tatoeba~\cite{tiedemann-2020-tatoeba}, and TICO-19~\cite{anastasopoulos-etal-2020-tico}. To summarize, our contributions are as follows:
\begin{itemize}[noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt]
\item We propose SMaLL-100, a shallow multilingual NMT model, focusing on low-resource language pairs.
\item We evaluate SMaLL-100 on several low-resource NMT benchmarks.
\item We show that our model significantly outperforms previous multilingual models of comparable size while being faster at inference. Additionally, it achieves comparable results with M2M-100~(1.2B) model, with 4.3$\times$ faster inference and a 3.6$\times$ smaller size.
\item While SMaLL-100 reaches 87.2\% performance of the 12B teacher model, we show that this gap can be closed with a few fine-tuning steps both for low and high-resource languages. 
\end{itemize}

\section{Model and Training}
\label{sec:method}
\newcommand{\Lagr}{\mathcal{L}}

\subsection{SMaLL-100 Architecture} 
It has been shown by \citet{kasai2021_deepshallow} that deep encoder / shallow decoder architectures can achieve good translation quality while being significantly faster at inference. \citet{berard-etal-2021-efficient} have confirmed that this is also valid for multilingual NMT. Here, we use a 12-layer Transformer encoder \cite{transformer} and 3-layer decoder. Table \ref{app:tab:hyper-params} in the Appendix \ref{app:impl} reports further details of the SMaLL-100 architecture. Different from M2M-100 model, we use language codes in the encoder side, as it is shown to perform better with shallow decoder architectures~\cite{berard-etal-2021-efficient}.  

\subsection{Training Strategy}
SMaLL-100 is trained with a combination of two loss functions: a standard Cross Entropy loss (CE) and a Knowledge Distillation loss (KD). Given source sequence $X$ and gold target translation $Y=(y_0,...,y_m)$, the CE loss is calculated as:
\begin{align}
\begin{split}
\Lagr_{ce} = - \sum_{j=0}^{m}\sum_{z=1}^{|K|} \mathbbm{1}\{y_j = z \} \operatorname{log}~p(y_j=z | y_{<j},X,\theta_S) 
\end{split}
\end{align}
where $|K|$ is the target vocabulary size, $\mathbbm{1}$ is the indicator function, and $\theta_S$ is the model parameters. $p()$ is the conditional probability function. \\
We additionally use a word-level distillation loss, which is the Kullback–Leibler divergence between the output distributions of the student and teacher models~\cite{hu-etal-2018-attention}. Specifically, it is calculated as:
\begin{align}
\begin{split}
\Lagr_{kd} = - \sum_{j=0}^{m}\sum_{z=1}^{|K|} q(y_j=z | y_{<j},X,\theta_T) \\
\times \operatorname{log}~p(y_j=z | y_{<j},X,\theta_S) 
\end{split}
\end{align}
where $\theta_T$ is parameters of the teacher model. $q()$ is the conditional probability of the teacher model. The total loss is computed as: 
\begin{align}
\begin{split}
\Lagr_{total} = \Lagr_{ce} + \alpha \Lagr_{kd}
\end{split}
\end{align}
where $\alpha$ is a trainable parameter. 

\subsection{Training Data} 
Our training data includes parallel sentences from CCMatrix~\cite{https://doi.org/10.48550/arxiv.1911.04944} and CCAligned~\cite{el-kishky-etal-2020-ccaligned} datasets, which are part of the training data used by \citet{m2m-100} to train the M2M-100 models. As our goal is to maintain the performance of low-resource languages, we balance the training data across all language pairs; specifically, 100K sentence pairs are sampled for each language pair.\footnote{For language pairs with less than 100K sentence pairs, we repeat their data. We randomly select 100K sentences for language pairs with more than 100K training sentences.} As a result, our training data contains nearly 456M parallel sentences, which is less than 6$\%$ of the original data on which M2M-100~\cite{m2m-100} was trained. We use the same languages as M2M-100.
\section{Experimental Setup}
\label{sec:setup}

\begin{table}
\centering
\begin{adjustbox}{width=0.7\linewidth}
\begin{tabular}{lc}

Resource Type & Criteria \\
\toprule
Very-Low & $|K| \leq 100K$  \\
Low & $100K < |K| \leq 1M $ \\
Medium & $1M < |K| \leq 100M $ \\
High & $100M < |K|$ \\
\bottomrule
\end{tabular}
\end{adjustbox}
\caption{\label{tab:langdist} The criteria to split languages into different resource categories. $|K|$ is the amount of training data to/from English.}
\end{table}

\begin{table*}
	\begin{adjustbox}{width=\textwidth}
	\centering
	\begin{tabular}{lcc|cccccccccccc|c}\toprule
		 &\multicolumn{15}{c}{\hspace{2cm} Language Direction} \\
		\cmidrule{4-15} 
		Model & params & Speed & VL2VL & VL2L & VL2M & VL2H & L2VL & L2L & L2M & L2H & M2VL & M2L & H2VL & H2L & AVG \\ \midrule
		\textbf{FLORES-101} & & & &  &  &  & &  &  &  &  & &  & &  \\[0.1ex] 
		FLORES-124 & 175M & 5.3$\times$ & 3.3 & 3.4 & 6.0 & 7.8 & 3.7 & 3.1 & 6.9 & 8.8 & 6.9 & 5.2 & 8.1 & 6.0 & 5.8\\
		M2M-100 & 418M & 3.1$\times$ & 4.3 & 3.7 & 7.8 & 9.4 & 5.4 & 3.4 & 9.1 & 11.3 & 9.9 & 5.8 & 11.4 & 6.6 & 7.3 \\
		FLORES-124 & 615M & 2.9$\times$ & 5.1 & 5.1 & 9.2 & 11.2 & 5.8 & 4.7 & 10.6 & 13.1 & 10.3 & 7.6 & 11.5 & 8.5 & 8.6 \\
	    \textit{Finetuned-100} & 330M & 7.8$\times$ & 6.1 & 5.4 & 8.7 & 11.3 & 5.7 & 4.1 & 9.0 & 11.8 & 10.4 & 6.8 & 13.0 & 8.0 & 8.4 \\
	    \textit{SMaLL-100} & 330M & 7.8$\times$ & \underline{7.9} & \underline{7.0} & 10.3 & 12.6 & 8.4 & \underline{6.1} & 11.6 & 14.3 & \underline{13.7} & \underline{9.0} & \underline{16.7} & \underline{10.2} & \underline{10.7} \\
	    \midrule
	    M2M-100 & 1.2B & 1.8$\times$ & 6.7 & 6.1 & \underline{10.8} & \underline{12.8} & \underline{8.7} & \underline{6.1} & \underline{13.0} & \underline{15.9} & 13.6 & 8.8 & 15.4 & 9.7 & 10.6 \\
	    M2M-100 & 12B & 1$\times$ & \textbf{8.7} & \textbf{8.8} & \textbf{11.9} & \textbf{13.7} & \textbf{11.7} & \textbf{9.7} & \textbf{15.4} & \textbf{18.2} & \textbf{16.5} & \textbf{12.6} & \textbf{18.7} & \textbf{13.9} & \textbf{13.3} \\
          \midrule \midrule
	   \textbf{Tatoeba} & & & &  &  &  & &  &  &  &  & &  & &  \\[0.1ex] 
		FLORES-124 & 175M  & 5.3$\times$ & - &7.6 & 15.7 & 10.1 & 4.6 & 5.3 & 11.5 & 10.8 & 14.0 & 10.2 & 6.4 & 7.5 & 9.4 \\
		M2M-100 & 418M  & 3.1$\times$ &- & 7.4 & 19.7 & 12.3 & 5.9 & 5.3 & 13.8 & 13.2 & 14.9 & 11.7  & 7.7 & 9.0 & 10.9 \\
		FLORES-124 & 615M  & 2.9$\times$ & - & \textbf{9.1} & 19.4 & 11.4 & 6.9 & \underline{7.6} & 12.7 & 13.7 & 14.4 & 13.3 & 8.0 & 9.7  & 11.4 \\
	    \textit{Finetuned-100} & 330M  & 7.8$\times$ & - & 4.0 & 21.1 & \underline{14.4} & 7.7 & 5.2 & 15.3 & 14.2 & 14.0 & 12.1 & 8.9 & 8.3 & 11.4 \\
	    \textit{SMaLL-100} & 330M  & 7.8$\times$ & - & 4.6 & \underline{22.1} & \textbf{16.4} & \underline{8.7} & 7.0 & \underline{16.7} & 15.8 & 16.3 & \underline{14.5} & \underline{10.6} & \underline{11.2} & \underline{13.1} \\
	    \midrule
	    M2M-100 & 1.2B  & 1.8$\times$ &- & \underline{8.8} & 19.5 & 13.1 & \underline{8.7} & 7.2 & 16.3 & \underline{17.0} & \underline{17.2} & 13.4 & \textbf{10.7} & 11.1  & 13.0 \\
	    M2M-100 & 12B & 1$\times$ & - & 8.6 & \textbf{23.5} & 13.1 & \textbf{9.8} & \textbf{10.2} & \textbf{17.8} & \textbf{17.9} & \textbf{18.5} & \textbf{15.2} & \textbf{10.7} & \textbf{13.2} & \textbf{14.4} \\ \midrule \midrule

	    \textbf{TICO-19} & & & &  &  &  & &  &  &  &  & &  & &  \\[0.1ex] 
		FLORES-124 & 175M & 5.3$\times$ & 4.6 & 5.5 & 8.1 & 11.5 & 4.4 & 5.6 & 9.7& 12.2& 3.9& 8.0& 4.2& 8.7 & 7.2\\
		M2M-100 & 418M & 3.1$\times$ & 4.0 & 5.5 & 9.8 & 13.7 & 4.2 & 5.7 & 11.6 & 14.9 & 4.1 & 8.8 & 5.3 & 9.4 & 8.1 \\
		FLORES-124 & 615M & 2.9$\times$ &  4.6 & 7.4 & 11.5 & 16.4 & 4.8 & 7.6 & 12.9 & 16.7 & 4.4 & 10.7 & 4.4 & 11.5 & 9.4 \\
	    \textit{Finetuned-100} & 330M & 7.8$\times$ & 6.1 & 7.2 & 11.9 & 17.4 & 5.5 & 6.1 & 12.1 & 15.2 & \underline{6.4} & 9.0 & \underline{9.5} & 10.3 & 9.7 \\
	    \textit{SMaLL-100} & 330M & 7.8$\times$& \textbf{7.8} & \underline{8.8} & \underline{13.3} & \underline{19.0} & \textbf{8.0} & 8.5 & \underline{14.3} & 17.8 & \textbf{8.3} & \underline{11.5} & \textbf{11.3} & \underline{12.7} & \underline{11.8} \\ 
	    \midrule
	    M2M-100 & 1.2B & 1.8$\times$ & 5.4 & 8.2 & 13.2 & 18.9 & 6.0 & \underline{8.7} & 14.0 & \underline{19.2} & 5.2& \underline{11.5} & 6.1& 12.5 & 10.8 \\
	    M2M-100 & 12B & 1$\times$ & \underline{6.4} & \textbf{10.9} & \textbf{15.4} & \textbf{20.6} & \underline{7.8} & \textbf{11.9} & \textbf{16.6} & \textbf{21.4} & \underline{6.4} & \textbf{15.4} & 8.7& \textbf{16.4} & \textbf{13.1} \\
	    \midrule
		\bottomrule
	\end{tabular}
	\end{adjustbox} 
	\caption{\label{table:main} Average spBLEU performance on FLORES-101, Tatoeba, and TICO-19 benchmarks for different language pair categories, defined in Appendix~\ref{app:detail-eval}. FLORES-101 results are computed on language pairs where M2M-100 12B has spBLEU scores higher than 3 to avoid polluting the analysis with meaningless scores. The first and second columns give the model size and speed-up ratios compared to M2M-100~(12B). Last column is the average spBLEU performance over all mentioned language directions. The best scores are shown in bold, and the second best results are shown with underline.}
\end{table*}

\subsection{Evaluation Benchmarks}

\paragraph{FLORES-101} is a multilingual NMT benchmark, containing 3,001 sentences from different domains, that are derived from English Wikipedia. Sentences are translated into 101 languages by human translators~\cite{flores101}. It mostly includes low and medium-resource languages. We use {\tt devtest} subset for the evaluation.

\paragraph{Tatoeba} is a crowd-sourced collection of user-provided translations in different languages~\cite{tiedemann-2020-tatoeba}. We choose a subset of languages from {\tt test} set of Tatoeba Challenge,\footnote{\url{https://github.com/Helsinki-NLP/Tatoeba-Challenge}} which are covered by M2M-100. 

\paragraph{TICO-19} was created during the COVID-19 pandemic~\cite{anastasopoulos-etal-2020-tico}. It contains sentences from 36 languages in the medical domain, including 26 low-resource languages. We evaluate on languages which are covered by M2M-100~\cite{m2m-100}. 

Inspired by \citet{flores101}, we split the languages based on the amount of available training sentences aligned with English into 4 different categories: Very-Low~(VL), Low~(L), Medium~(M), and High-resource~(H). As the true amount of training data is both dependent on quality and quantity of parallel sentences, \citet{flores101} suggested to estimate it by computing the number of bitext data aligned with English, that is calculated from statistics of OPUS corpora~\cite{tiedemann-2012-parallel}. Table~\ref{tab:langdist} illustrates the criteria for choosing the category of different languages. More details about the distribution of language pair categories in each benchmark are provided in Appendix~\ref{app:detail-eval}. 

\subsection{Baselines}

\paragraph{M2M-100}\citet{m2m-100} is a recent many-to-many NMT model covering 100 languages. \citet{m2m-100} provide 3 variants with respectively 418M, 1.2B, and 12B parameters. We compare against these 3 variants.
\paragraph{FLORES-124} is an extension of M2M-100, covering additional 24 languages. Training data of the additional languages is derived from OPUS~\cite{tiedemann-2012-parallel}. \citet{flores101} provide two models with 175M and 615M parameters. We use both models as baselines. 
\paragraph{FineTuned-100} uses the same architecture as defined in Section~\ref{sec:method}, but KD loss~($\Lagr_{kd}$) is not used for training. For a fair comparison, it is trained for the same number of steps as SMaLL-100 model. 

\begin{table*}
\centering
\begin{adjustbox}{width=0.9\linewidth}
\begin{tabular}{lccccc}\toprule
    & &\multicolumn{3}{c}{Fine-tuned} \\
    \cmidrule{3-5} 
Language pair & Language Type & M2M-100 (418M) & SMaLL-100 & steps & M2M-100 (12B) \\ \midrule
Cebuano-English & Low & 18.6 & \textbf{29.8} & 0.5K & 27.7   \\
English-Igbo & Low & 9.2 & \textbf{15.7}  & 1.5K & 14.9 \\
English-Malayalam & Low & 16.7 & \textbf{20.8}  & 3K & 20.6 \\
Georgian-Russian & Medium & 6.9 & \textbf{13.1} & 0.5K & 10.1 \\
English-Italian & High & 33.3 & 33.4 & 20K & \textbf{33.5} \\
French-Italian & High & 31.3 & 31.5 & 20K & \textbf{32.0} \\
Italian-Spanish & High & 26.6 & 26.8 & 20K & \textbf{27.0} \\
\bottomrule
\end{tabular}
\end{adjustbox}
\caption{\label{app:fine-tune} spBLEU performance of fine-tuned SMaLL-100 and M2M-100 (418M) for the specified step, and M2M-100~(12B) on FLORES-101 devtest. The "step" column is the number of training steps required to reach M2M-100~(12B) performance. The type of each language pair is defined as the minimum of source and target language categories.}
\end{table*}

\subsection{Implementation Details}

SMaLL-100 contains nearly 330M parameters with 12 encoder and 3 decoder Transformer layers.\footnote{It is initialized with M2M-100~(418M), using its first 3 decoder layers for the initialization of the student's decoder.} It is trained for 30 days on 16 TESLA V100-32GB GPUs,\footnote{4 GPUs are used for the training of the student model. 12 GPUs are utilized to do the model parallelism of the teacher model.} with a batch size of 1K tokens and accumulated gradients over 9 batches. We implement our model using fairseq repository.\footnote{\url{https://github.com/facebookresearch/fairseq}} We use {\tt last-checkpoint}\footnote{\url{https://github.com/facebookresearch/fairseq/tree/main/examples/m2m_100}} of M2M-100~(12B) for the teacher model. For decoding, the beam search of 5 is applied. All hyper-parameters regarding the architecture and optimization strategy are provided in Appendix~\ref{app:impl}. \\ 
For a faster convergence, we first fine-tune SMaLL-100 for 150k steps without distillation~($\Lagr_{kd}$). Then, it is trained with both losses for 756K steps~(nearly 1 epoch). For evaluation, we use SentencePiece BLEU~(spBLEU), as it is shown to be a fair metric in multilingual settings~\cite{flores101}.\footnote{It utilizes a SentencePiece tokenizer with 256K tokens:~\url{https://github.com/facebookresearch/flores}} We use the same tokenizer and dictionary as M2M-100.

\section{Results and Discussion}
\label{sec:result}

\subsection{Low-Resource NMT Benchmarks}

Table~\ref{table:main} shows the average spBLEU performance on FLORES-101, Tatoeba, and TICO-19 test sets for different categories of language directions.\footnote{Complete spBLEU calculations of different language pairs on tested NMT benchmarks are provided in Appendix~\ref{app:spbleu}. Speed is calculated on 2 TESLA V100-32GB GPUs with a batch size of 1 sentence over a subset of FLORES-101 devtest, containing nearly 10K sentences from all language pairs.} SMaLL-100 outperforms all the models with comparable sizes while being smaller and faster at inference. Specifically, it outperforms M2M-100 418M both in terms of performance (+3.1 spBLEU) and inference speed (2.5$\times$ faster). We believe that Finetuned-100 outperforms M2M-100 418M for low-resource languages thanks to finetuning on the balanced dataset.  The higher performance of SMaLL-100 compared to Finetuned-100 across all benchmarks shows the benefit of KD loss which allows to distill knowledge from the teacher model.  Additionally, SMaLL-100 achieves competitive results with M2M-100~(1.2B), while being 3.6$\times$ smaller and 4.3$\times$ faster at inference. Compared to the biggest M2M-100 model~(12B), SMaLL-100 loses nearly 1.7 spBLEU but is 36$\times$ smaller and 7.8$\times$ faster. Regarding medium and high-resource language pairs~(as shown in Appendix~\ref{app:spbleu:remain}), SMaLL-100 achieves better or similar performance compared to M2M-100~(418M) and FLORES-124~(615M), while it contains fewer parameters and is faster at the evaluation time. It under-performs for some medium and high-resource language pairs compared to the teacher model~(M2M-100 12B), which could be easily recovered, as we describe in the remaining section.

\subsection{Recovering Teacher Model Performance}
To go further, we demonstrate that SMaLL-100 can easily recover the performance of the teacher model with just a few fine-tuning steps, both for low and high-resource language pairs. For comparison, we fine-tune M2M-100 (418M) model with the same number of steps. \\
Table~\ref{app:fine-tune} reports spBLEU performance for several language pairs, alongside the number of fine-tuning steps, required by SMaLL-100 model to reach M2M-100 (12B) performance.\footnote{More dataset and implementation details of these fine-tuning experiments are provided in Appendix~\ref{app:detail-low}.}
We see that SMaLL-100 achieves better performance than M2M-100~(12B) after a few training steps on low-resource language pairs. For high-resource language pairs, SMaLL-100 is fine-tuned for 20K steps to reach the performance of M2M-100~(12B) model. Additionally, fine-tuned SMaLL-100 significantly outperforms fine-tuned M2M-100 (418M) model on low and medium-resource languages. This confirms that SMaLL-100 could be a powerful and lightweight initialization model for training on different language pairs.

\section{Related Work}
\label{relatedwork}

\paragraph{Compression and Distillation.} Over the past few years, pre-trained models lead to significant improvement by increasing the parameter size~\cite{raffel2019exploring,m2m-100,zhang2022opt}, which makes it challenging to use them in the resource-constraint environment. Previous work use several compression techniques e.g. knowledge distillation~\cite{kim-rush-2016-sequence,light2021}, pruning~\cite{behnke-heafield-2020-losing,zhang-etal-2021-enlivening,mohammadshahi2022compressed}, and quantization~\cite{tao2022compression,yao2022zeroquant} to provide a reasonable-size model, while keeping the performance. 
\paragraph{Multilingual NMT.} It provides a single model to translate between any pair of languages, which significantly improves performance on low-resource languages thanks to knowledge transfer~\cite{haddow2021survey}. 
Several works~\cite{dong-etal-2015-multi,firat-etal-2016-multi,platanios-etal-2018-contextual,m2m-100,berard-etal-2021-efficient} propose to include both language-specific, and language-independent parameters in MNMT models. Recently, massively MNMT models~\cite{neubig-hu-2018-rapid,https://doi.org/10.48550/arxiv.1907.05019,aharoni-etal-2019-massively,m2m-100,zhang-etal-2020-improving} have been proposed to translate between more than 100 languages. However, these models usually contain a huge number of parameters to maintain performance in both high and low-resource languages. Different from the previous work, we introduce SMaLL-100, which outperforms previous models with comparable size in low-resource language directions, while achieving better speed and being smaller.

\section{Conclusion}

We presented SMaLL-100 model, a shallow multilingual NMT model, focusing on low-resource languages. We evaluated our model on different NMT benchmarks.
SMaLL-100 significantly outperforms multilingual models of comparable size on all of the tested benchmarks (FLORES-101, Tatoeba, TICO-19) and is much faster at inference.
It also achieves competitive results with M2M-100 1.2B~\cite{m2m-100}, while being 4.3$\times$ faster at inference and 3.6$\times$ smaller. Compared to M2M-100~(12B), the biggest available MNMT model, SMaLL-100 loses nearly 1.7 spBLEU on average but it is significantly faster~(7.8$\times$) and smaller~(36$\times$), which makes it a good fit for resource-constrained settings. Additionally, we show that SMaLL-100 can achieve similar performance as M2M-100~(12B) with just a few steps of fine-tuning on specific language pairs.
\section*{Limitations}

As mentioned in Section~\ref{sec:result}, SMaLL-100 model under-performs for some medium and high-resource languages, which could be resolved by further fine-tuning. Due to our computation constraint, we train SMaLL-100 model on nearly 6$\%$ of the original M2M-100 model. So, we encourage future research to increase the size of training data~(especially for low-resource languages) to achieve better performance. Also, future research could apply different distillation strategies~\cite{wu-etal-2020-skip,wang-etal-2021-selective}, as we just used word-level knowledge distillation loss~\cite{hu-etal-2018-attention}.
\section*{Acknowledgement}

The work is done during the research internship at NAVER LABS Europe. Alireza Mohammadshahi is supported by the Swiss National Science Foundation~(grant number CRSII5-180320).

\bibliographystyle{acl_natbib}
\bibliography{emnlp2021}

\newpage
\renewcommand\thesection{\Alph{section}}
\renewcommand\thesubsection{\thesection.\Alph{subsection}}
\setcounter{section}{0}
\onecolumn
\begin{appendices}

\section{Details of Evaluation benchmarks}
\label{app:detail-eval}

\subsection{Resource-type of Languages in Evaluated Datasets}
\label{app:lang-type}

\begin{table}[!ht]
\centering
\begin{adjustbox}{width=0.7\linewidth}
\begin{tabular}{cc|cc|cc|cc|cc}\toprule
af & Low & lg & Very-Low & lt & Medium & sn & Low & gl & Medium \\ \midrule
am & Low & ka & Medium & luo & Low & sd & Very-Low & gd & Low \\ \midrule
ar & Medium & de & High & lb & Medium & sk & Medium & ht & Low \\ \midrule
hy & Low & el & Medium & mk & Medium & sl & Medium & su & Low \\ \midrule
as & Very-Low & gu & Low & ms & Low & so & Low & ln & Very-Low \\ \midrule
ast & Low & ha & Low & ml & Low & ku & Low & ilo & Low \\ \midrule
az & Low & he & Medium & mt & Medium & es & High & mg & Medium \\ \midrule
be & Very-Low & hi & Medium & mr & Low & sw & Low & tn & Very-Low \\ \midrule
bn & Medium & hu & Medium & mi & Low & sv & Medium & br & Medium \\ \midrule
bs & Low & is & Medium & mn & Low & tg & Low & ns & Very-Low \\ \midrule
bg & Medium & ig & Low & ne & Very-Low & ta & Low & si & Medium \\ \midrule
my & Low & id & Medium & nso & Very-Low & te & Low & yi & Low \\ \midrule
ca & Medium & ga & Low & no & Medium & th & Medium & fy & Medium \\ \midrule
ceb & Low & it & High & ny & Low & tr & Medium & sq & Medium \\ \midrule
zh & Medium & ja & Medium & oc & Very-Low & uk & Medium & ss & Very-Low \\ \midrule
hr & Very-Low & jv & Medium & or & Very-Low & umb & Low & fr & High \\ \midrule
cs & Medium & kea & Very-Low & om & Low & ur & Low & ff & Very-Low \\ \midrule
da & Medium & kam & Very-Low & ps & Low & uz & Very-Low & lo & Low \\ \midrule
nl & Medium & kn & Low & fa & Medium & vi & Medium & lv & Medium \\ \midrule
en & High & kk & Low & pl & Medium & cy & Low & ru & High \\ \midrule
et & Medium & km & Low & pt & High & wo & Very-Low & sr & Medium \\ \midrule
tl & Very-Low & ko & Medium & pa & Low & xh & Low & zu & Low \\ \midrule
fi & Medium & ky & Low & ro & Medium & yo & Low & ba & Low \\ 
\bottomrule
\end{tabular}
\end{adjustbox}
\caption{\label{tab:langtype} ISO-639 code and resource type of languages used in evaluated NMT benchmarks.}
\end{table}

\subsection{FLORES-101}

We use {\tt devtest} subset of FLORES-101 for the evaluation. To better compare different models, we exclude evaluation of language pairs, in which the spBLEU performance of M2M-100 12B~\cite{m2m-100} model is below 3. This gives 5,934 language directions for the comparison. Table~\ref{app:tab:flores-dist} shows the distribution of different categories of language pairs.

\begin{table}[!ht]
\centering
\begin{adjustbox}{width=\linewidth}
\begin{tabular}{lcccccccccccccccc}\toprule
 & VL2VL & VL2L & VL2M & VL2H & L2VL & L2L & L2M & L2H & M2VL & M2L & M2M & M2H & H2VL & H2L & H2M & H2H \\ \midrule
No. lang. pairs & 44 & 200 & 388 & 81 & 144 & 645 & 960 & 186 & 181 & 992 & 1330 & 259 & 35 & 190 & 257 & 42 \\
\bottomrule
\end{tabular}
\end{adjustbox}
\caption{\label{app:tab:flores-dist} Distribution of resource categories for different language directions on FLORES-101~\cite{flores101}.}
\end{table}

\subsection{Tatoeba Challenge}
\label{app:tatoeba:data}

We use the {\tt test} subset data, provided by \citet{tiedemann-2020-tatoeba}\footnote{\url{https://github.com/Helsinki-NLP/Tatoeba-Challenge}} to evaluate all models. We choose a subset of dataset that includes languages which are covered by M2M-100~\cite{m2m-100} model. This brings 1,844 language pairs for the evaluation. The distribution of different language pair categories is shown in Table~\ref{app:tab:tatoeba-dist}. 

\begin{table}[!ht]
\centering
\begin{adjustbox}{width=\linewidth}
\begin{tabular}{lccccccccccccccc}\toprule
 & VL2L & VL2M & VL2H & L2VL & L2L & L2M & L2H & M2VL & M2L & M2M & M2H & H2VL & H2L & H2M & H2H \\ \midrule
No. lang. pairs &  7 & 30 & 37 & 7 & 34 & 144 & 113 & 30 & 144 & 632 & 237 & 37 & 113 & 237 & 42\\
\bottomrule
\end{tabular}
\end{adjustbox}
\caption{\label{app:tab:tatoeba-dist} Distribution of resource categories for different language directions on Tatoeba Challenge.}
\end{table}

\subsection{TICO-19}

We use the evaluation benchmark provided by \citet{anastasopoulos-etal-2020-tico}\footnote{\url{https://tico-19.github.io/}} to compare all models in multilingual medical domain. We utilize language pairs that are included in M2M-100~\cite{m2m-100} model. This gives us 650 language pairs for the evaluation. Table~\ref{app:tab:tico19-dist} shows the number of language pairs in different resource types of language directions.

\begin{table}[!ht]
\centering
\begin{adjustbox}{width=\linewidth}
\begin{tabular}{lcccccccccccccccc}\toprule
 & VL2VL & VL2L & VL2M & VL2H & L2VL & L2L & L2M & L2H & M2VL & M2L & M2M & M2H & H2VL & H2L & H2M & H2H \\ \midrule
No. lang. pairs &  12 & 48 & 24 & 16 & 48 & 132 & 72 & 48 & 24 & 72 & 30 & 24 & 16 & 48 & 24 & 12 \\
\bottomrule
\end{tabular}
\end{adjustbox}
\caption{\label{app:tab:tico19-dist} Number of language pairs for different categories of language directions on TICO-19.}
\end{table}

\section{Hyper-Parameters for Architecture and Optimization}
\label{app:impl}

\begin{table}[!ht]
\centering
\begin{adjustbox}{width=0.8\linewidth}
\begin{tabular}{lc|lc}\toprule
Hyper-Parameter & Specification & Hyper-Parameter & Specification\\ \midrule
Encoder Layer & 12 & Scheduler & inverse-sqrt\\
Decoder Layer & 3 & Optimizer & Adam \\
Encoder Emb dim. & 1024 & Clip Norm & 1.0 \\
Decoder Emb dim & 1024 & Learning Rate & 1e-4 \\
Encoder FFN Emb dim. & 4096 & Warmup init LR & 1e-07 \\
Decoder FFN Emb dim. & 4096 & Warmup Updates & 40K \\
Number of attn heads & 16 & Adam Betas & 0.9,0.98\\
Attention Dropout & 0.1 & Adam eps & 1e-6 \\
Share Encoder/Decoder Emb. & True & Label Smoothing & 0.1 \\
FP16 & True & Dropout & 0.1 \\
Loss Scalar & 2.0 & Max Tokens (per GPU) & 1,000 \\

\bottomrule
\end{tabular}
\end{adjustbox}
\caption{\label{app:tab:hyper-params} List of hyper-parameters used for the architecture, and optimization. }
\end{table}

\newpage 
\section{spBLEU Results}
\label{app:spbleu}
~~
\subsection{spBLEU of Remaining Language Directions}
\label{app:spbleu:remain}

\begin{table}[!ht]
\centering
	\begin{adjustbox}{width=0.6\textwidth}
	\centering
	\begin{tabular}{lcccccc}\toprule
		 &\multicolumn{5}{c}{\hspace{3.5cm} Language Direction} \\
		\cmidrule{4-7} 
		Model & params & Speed & M2M & M2H & H2M & H2H \\ \midrule
		\textbf{FLORES-101} \\[1ex] 
		FLORES-124 & 175M & 5.3$\times$ & 13.7 & 18.0 & 16.8 & 23.0 \\
		M2M-100 & 418M & 3.1$\times$ & 18.1 & 23.0 & 21.6 & 27.8 \\
		FLORES-124 & 615M & 2.9$\times$ & 19.3 & 24.1 & 22.7 & 28.9 \\
	    \textit{Finetuned-100} & 330M & 7.8$\times$ & 16.4 & 21.2 & 19.7 & 26.0 \\
	    \textit{SMaLL-100} & 330M & 7.8$\times$ & 19.3 & 24.2 & 22.6 & 28.8 \\
	    
	    \midrule
	    M2M-100 & 1.2B & 1.8$\times$ & 22.3 & 28 & 25.8 & 32.7 \\
	    M2M-100 & 12B & 1$\times$ & 23.9 & 29.5 & 27.6 & 34.3 \\		\midrule \midrule

	   \textbf{Tatoeba} \\[1ex] 
		FLORES-124 & 175M  & 5.3$\times$ & 25.2 & 28.1 & 24.5 & 35.3 \\
		M2M-100 & 418M  & 3.1$\times$ & 29.6 & 34.0 & 29.6 & 41.6 \\
		FLORES-124 & 615M  & 2.9$\times$ & 31.7 & 35.5 & 31.0 & 43.0 \\
	    \textit{Finetuned-100} & 330M  & 7.8$\times$ & 28.3 & 34.2 & 28.1 & 39.2 \\
	    \textit{SMaLL-100} & 330M  & 7.8$\times$ & 31.9 & 36.3 & 31.1 & 42.4 \\
	    \midrule
	    M2M-100 & 1.2B  & 1.8$\times$ & 33.8 & 39.0 & 34.2 & 47.4 \\
	    M2M-100 & 12B & 1$\times$ & 33.1 & 39.0 & 34.2 & 48.7 \\ \midrule \midrule

	    \textbf{TICO19}  \\[1ex] 
		FLORES-124 & 175M & 5.3$\times$ & 15.1 & 20.3 & 17.7 & 28.1 \\
		M2M-100 & 418M & 3.1$\times$ & 20.6 & 26.6 & 23.6 & 33.1 \\
		FLORES-124 & 615M & 2.9$\times$ &  20.2 & 27.0 & 23.3 & 33.9 \\
	    \textit{Finetuned-100} & 330M & 7.8$\times$ & 19.7 & 24.4 & 23.4 & 31 \\
	    \textit{SMaLL-100} & 330M & 7.8$\times$& 21.7 & 27.4 & 25.2 & 33.7 \\ 
	    \midrule
	    M2M-100 & 1.2B & 1.8$\times$ & 21.1 & 30.2 & 24.8 & 37.7 \\
	    M2M-100 & 12B & 1$\times$ & 24.8 & 33.1 & 28.3 & 39.4 \\
	    \midrule
		\bottomrule
	\end{tabular}
	\end{adjustbox} 
	\caption{ Average spBLEU performance of different models on FLORES-101, Tatoeba, and TICO-19 benchmarks for medium and high-resource language pair categories, defined in Section~\ref{sec:result}. The FLORES-101 results are computed on language pairs where M2M-100 12B has spBLEU scores higher than 3 to avoid polluting the analysis with meaningless scores. The first and second columns give the model size and speed-up ratios compared to M2M-100~(12B).}
\end{table}

\newpage
\subsection{FLORES-101}

﻿\begin{table}[!ht]
    \centering
    \begin{adjustbox}{width=\linewidth}

\end{adjustbox}
\caption{spBLEU performance of last checkpoint of SMaLL-100 model on language pairs of FLORES-101.}
\end{table}

\subsection{Tatoeba}

﻿\begin{table}[!ht]
    \centering
    \begin{adjustbox}{width=\linewidth}
    %
\end{adjustbox}
\caption{spBLEU performance of last checkpoint of SMaLL-100 model on language pairs of Tatoeba.}
\end{table}

\newpage
\subsection{TICO19}

﻿\begin{table}[!ht]
    \centering
    \begin{adjustbox}{width=\linewidth}
    %

\end{adjustbox}
\caption{spBLEU performance of last checkpoint of SMaLL-100 model on language pairs of TICO19.}
\end{table}

\section{Details of Fine-Tuning on Low-Resource Language Pairs}
\label{app:detail-low}

For further fine-tuning of SMaLL-100 and M2M-100~(418M) models on selected language pairs, we use bilingual data provided by \citet{tiedemann-2020-tatoeba}~(release 2021.08.07)\footnote{\url{https://github.com/Helsinki-NLP/Tatoeba-Challenge/blob/master/data/README-v2021-08-07.md}} as its training data is less noisy. We evaluate fine-tuned models on {\tt devtest} subset of FLORES-101~\cite{flores101} benchmark with spBLEU metric~\cite{flores101}. We use the same hyper-parameters, as defined in Appendix~\ref{app:impl}. We train each model on 2 TESLA V100-32GB GPUs.

\end{appendices}

\end{document}